OpenStudy (anonymous):

ad/c+ae/b=3(-15)/-5+3(4)/12=?????

OpenStudy (ybarrap):

$$ \tt \Large\Large~~ad/c+ae/b\\ \Large=\frac{-3\times(-15)}{-5}+\frac{3\times 4}{12}\\ \Large=-3\times (3)+\frac{3\times 4}{3\times 4}\\ \Large=-9+1\\ \Large=-8 $$ Does this make sense?

OpenStudy (ybarrap):

Your're trying to solve this, right? $$ =\frac{-3\times(-15)}{-5}+\frac{3\times 4}{12}\\ $$

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the numbers I gave go where the letters are in the problem I gave

OpenStudy (anonymous):

you replace letters w/numbers.. your format of how ur wording things I completely don't understand

OpenStudy (ybarrap):

So a=3, b=12, c=-5, d=-15, and e=4? And you're trying to solve this equation? $$ =\frac{3\times(-15)}{-5}+\frac{3\times 4}{12}\\ $$

OpenStudy (ybarrap):

Do you agree that these are the same? $$ \tt \Large~~ad/c+ae/b\\ \Large=\frac{a\times d}{c}+\frac{a\times e}{b}\\ $$

OpenStudy (anonymous):

don't understand ur format sorry

OpenStudy (ybarrap):

I am writing d/c as \( \dfrac{d}{c}\) This is a fairly common way too write a fraction. So, for example, 3/2 is the same as \( \dfrac{3}{2}\) Still doesn't make sense?

OpenStudy (anonymous):

no

OpenStudy (ybarrap):

ok, let's keep your format ad/c+ae/b =3(-15)/-5+3(4)/12, but (-15)/-5 = 3 So this becomes 3(3) +3(4)/12 = 9 + 3(4)/12 So far ok?

OpenStudy (anonymous):

no. I got this next. 3(-15)/-5+3(4)/12= -45/-5+12/12= 540/60

OpenStudy (anonymous):

you don't distribut like terms. this is order of operations with rational numbers.. alg 1 homework

OpenStudy (anonymous):

I have 10 questions if u don't care to stay on the line helping with all ten pleaseee?????

OpenStudy (ybarrap):

Close. -45/-5 + 12/12 = -45(-12)/60 + 60/60 =(-45(-12)+60)/60 =(540 + 60)/60 =600/60 =10

OpenStudy (ybarrap):

Does this make sense?

OpenStudy (anonymous):

whatd u do in the 2nd line????

OpenStudy (anonymous):

yeah

OpenStudy (ybarrap):

I'm finding a common denominator, which is the common denominator between -5 and 12, which is 60.

OpenStudy (anonymous):

oh ok yeah it makes sense, I have to copy that down real quick. then could u help with more please I have 9 questions more. if u help me with all ten ill give a best response

OpenStudy (ybarrap):

I would have done it a little different for that line. I would have done like this: -45/-5 + 12/12 = -45/-5 + -5/-5 =(-45-5)/-5 =(-50)/-5 =10 Working with -5 is easier than that big number 60

OpenStudy (anonymous):

tooi still understand it lol (: thanks. next question?

OpenStudy (ybarrap):

yep

OpenStudy (anonymous):

bc/d-be/a=12(-5)/-15-12(4)/3 common denominator -15

OpenStudy (anonymous):

you there?

OpenStudy (ybarrap):

just got back

OpenStudy (ybarrap):

yes, common denominator is -15.Good.

OpenStudy (anonymous):

how do u solve it????? lol

OpenStudy (ybarrap):

ok so you have, =12(-5)/-15-12(4)/3 =12(-5)/-15-12(4)(-5)/-15 Right?

OpenStudy (anonymous):

yeahhh

OpenStudy (ybarrap):

But let's make this easier. Let's simplify before finding a common denominator: =12(-5)/-15-12(4)/3 =12/3-4(4) Do you see why?

OpenStudy (anonymous):

no

OpenStudy (ybarrap):

=12(-5)/-15-12(4)/3 For the first term, -5/-15 = 1/3 and for the second term, 12/3 = 4

OpenStudy (anonymous):

ok?????

OpenStudy (ybarrap):

So, instead of 12(-5)/-15, we can write 12/3 Instead of 12(4)/3, we can write 4(4) Combine this and we get 12/3 - 4(4) = 4-16=-12

OpenStudy (ybarrap):

No need for finding denominators now

OpenStudy (anonymous):

no I don't think u can take 5/15 12 times 5 =? divided by 15 you lost me on first step

OpenStudy (anonymous):

i don't think your doing it right. no offense but i think you might be misinterpreting my questions

OpenStudy (anonymous):

i could take a pic and post???

OpenStudy (ybarrap):

12(-5)/-15 = 12 times -5/-15 = 12 times 1/3 = 12 times 1/3 = 12/3 = 4 You know that "ab/c" is the same as "a times b/c right"? Go ahead and post it.

OpenStudy (anonymous):

camera keeps blurring cant. you really aren't making sense. maybe you could draw steps and answer??

OpenStudy (anonymous):

with fraction bar --- instead of /

OpenStudy (ybarrap):

12(-5) 12 ______ = ___ = 4 -15 3 Is this better?

OpenStudy (anonymous):

yes ! (:

OpenStudy (ybarrap):

But does it make sense?

OpenStudy (anonymous):

yeah but u messed up on writing the problem

OpenStudy (ybarrap):

ok, Tell me the correct problem

OpenStudy (ybarrap):